OpenStudy (anonymous):

x^2 + 3x = -25 solve by completing the square

OpenStudy (anonymous):

x^2+3x+25=0 (x+1.5)^2+23.75=0 (x+1.5)^2=-23.75 this equation has no roots

OpenStudy (anonymous):

okay, thank you